OpenStudy (anonymous):

derivative of y =cot^2(sinx)

OpenStudy (dumbcow):

use chain rule u = sin x du = cos x now find \[\frac{d}{du} \cot^2 u\] \[=2 \cot u (-\csc^2 u)\] plug u back in and multiply by du \[\rightarrow -2 \csc^2 (\sin x) \cot (\sin x) \cos x\]